OpenStudy (anonymous):

(2222^5555+5555^2222) what is the remainder when divided by 7 ?

OpenStudy (anonymous):

Did you learn module before?

OpenStudy (anonymous):

no ,what is that ? can u guide me?

OpenStudy (anonymous):

@hartnn PLZ SEE THIS

hartnn (hartnn):

2222 = 2*1111 = 2* 11 * 101 5555= 5*11*101 101 can be factored further ?

OpenStudy (anonymous):

what will be the answer plz show the further steps ?

OpenStudy (anonymous):

It is more easy to do with module. I can't see any simple solution to this question.

OpenStudy (lgbasallote):

there is a simple solution actually... 2^1 = 2 2^2 = 4 2^3 = 8 2^4 = last digit 6 2^5 last digit 2 so you see... 2 repeats the loop for every exponent of 4 you have 2222^5555 no matter how big the answer is...since the last digit is 2 and the exponent is divisible by 5... the last digit of 2222^5555 will be 4 then... 5^1 = 5 5^2 = 5 5^3 = 5 as you can see...all powers of 5 have 5 as the last digit.. so...5555^2222..no matter how big this is...the last digit is just 5 do these make sense?

OpenStudy (anonymous):

Form the pattern as Lg told...

OpenStudy (shubhamsrg):

well.. 2222 = 3 mod 7 =>2222^5555 = 3^5555 mod7 and 5555 = 4 mod 7 =>5555^2222 = 4^2222 mod7 now,, 3^3 = -1 mod7 (3^3)^(1851) = (-1)^(1951) mod 7 3^5553 = (-1) mod 7 =>3^(5555) = -9 mod(7) = 2 mod(7) same way,,we can see,, 4^3 = 1 mod 7 (4^3)^(740) = 1^740 mod7 4^2220 = 1 mod7 =>4^2222 = 16mod 7 = 2 mod 7 so final remainder will be 2+2 = 4..

OpenStudy (anonymous):

\[\frac{ 10+5 }{ 7 } = \frac{ 10 }{ 7 } + \frac{ 5 }{ 7 }\] Combine this concept and @lgbasallote ' s method...)

OpenStudy (shubhamsrg):

@igba last digit of 2222^5555 should be 8 and not 4 anyways,,what use will you make of that last digit thing anyways ?

OpenStudy (shubhamsrg):

am not arguing,just curious..

OpenStudy (anonymous):

i mean the pattern (remainder) 2^1 /7 = 2 2^2 /7 = 4 2^3 /7 = 1 2^4 /7 = 2 etc....this pattern continues.......adn Find Such pattern for 5^1/7 ......

OpenStudy (lgbasallote):

let's let wolfram judge @shubhamsrg

OpenStudy (lgbasallote):

http://www.wolframalpha.com/input/?i=+2222%5E5555 seems 4 is the last digit

OpenStudy (anonymous):

I do not think premkishan could understand what you guys said.

OpenStudy (shubhamsrg):

i think theres' a ... there after "4" + a option of "more digits" anyways, i dont need wolfram to crosscheck..its easy to see.. 2^(4n+1) has last digit 2 2^(4n+2) has last digit 4 and 4n+3 -->8 5555 = 4n+3

OpenStudy (lgbasallote):

@NewbieCarrot how sure are you he cares?

OpenStudy (shubhamsrg):

atleast i didnt tell anything wrong ! o.O

OpenStudy (shubhamsrg):

he has to learn modulus for such kinda questions,,as simple as that..

OpenStudy (anonymous):

I think there are some reason why he have this question without learning module.

OpenStudy (anonymous):

Or he was not telling the truth.

OpenStudy (lgbasallote):

if you say you don't need wolfram to solve these....then can you tell me why you did not get the same result for the remainder as wolfram did? http://www.wolframalpha.com/input/?i=%282222%5E5555%2B5555%5E2222%29+mod+7

OpenStudy (shubhamsrg):

ohh..lol..leme recheck..

OpenStudy (anonymous):

-9 mod(7) = 5 mod(7) not 3 (mod 7)

OpenStudy (shubhamsrg):

yes,,right !! silly of me !! hmm..

OpenStudy (shubhamsrg):

so with that correction,,ans indeed becomes 0 .. long live moduli ! :P